OpenStudy (anonymous):

HELPPPPPP A person can order a new car with a choice of 10 possible colors, with or without air conditioning, with or without automatic transmission, with or without power windows, and with or without a CD player. In how many different ways can a new car be ordered with regard to these options?

OpenStudy (anonymous):

\[ 10\cdot 2\cdot 2\cdot 2\cdot 2 \]

OpenStudy (anonymous):

Since each choice is in a series, you multiply the number of options for each choice.

OpenStudy (anonymous):

thank you!